A box contains 4 blue balls 5 green balls and 3 yellow balls. You randomly select 3 balls, one at a time, from the box without replacing a selected ball. What is the probability of selecting a blue, a green, a yellow in that order?
Answer by stanbon(75887)

A box contains 4 blue balls 5 green balls and 3 yellow balls. You randomly select 3 balls, one at a time, from the box without replacing a selected ball. What is the probability of selecting a blue, a green, a yellow in that order?
-------------
P(blue)*P(green)*P(yellow)
= 4/12 * 5*11 * 3/10
= 5/110
= 1/22
===============
Cheers,
Stan H.